4/(x-4)^2=1
i having alot of trouble with this one. I think i need to get the numerator over to the other side.

Answer by checkley71(8403)

YOU'RE RIGHT. CROSS MULTIPLY THE DENOMINATOR BY 1 AND THAT WILL BE SET = 4*1
4/(X-4)~2=1 OR 4/(X~2-8X+16)=1 OR X~2-8X+16=4 OR X~2-8X+12=0 OR (X-6)(X-2)=0 OR
X=6 & X=2